# L3Ms — Lagrange Large Language Models

### v1

@@ Abstract
Supervised fine-tuning (SFT) and alignment of large language models (LLMs) are key steps in providing a good user experience. However, the concept of an appropriate alignment is inherently application-dependent, and current methods often rely on heuristic choices to drive the optimization. In this work, we formulate SFT and alignment as a constrained optimization problem, where the LLM is trained on a task while being required to meet application-specific requirements, without resorting to heuristics. To solve this, we propose Lagrange Large Language Models (L3Ms), which employ logarithmic barriers to enforce the constraints. This approach allows for the customization of L3Ms across diverse applications while avoiding heuristic-driven processes. We demonstrate experimentally the versatility and efficacy of L3Ms in achieving tailored alignments for various applications.
@@ 1 Introduction
Large language models (LLMs) are used for a wide range of tasks: as chatbots (Brown et al., 2020; OpenAI, 2024), for code generation (Ahmad et al., 2021; Wang et al., 2021; Rozière et al., 2024), for medical assistance (Yang et al., 2022; Moor et al., 2023), and more. The key ingredients for their impressive downstream performance are supervised fine-tuning (SFT) and alignment; the former fine-tunes the LLM to a task of interest, while the latter instills it with preferential properties. Arguably, the right combination of preferential properties is highly application/task-dependent. For instance, a scholar might want a chatbot to be honest and factual to assist with their work, whereas a fiction writer might prefer the opposite behavior to help create fantastical imaginary worlds. There is also plenty of (anecdotal) evidence in support: some LLMs refuse to provide information on how to “kill” a process in Unix, recommending the use of less violent strategies for dealing with wayward computer programs instead. [^1] https://www.reddit.com/r/LocalLLaMA/comments/180p17f/new_claude_21_refuses_to_kill_a_python_process/ Therefore, we need frameworks for the customization of LLMs.
@@
Consequently, Li et al. (2021); Bai et al. (2022); Rame et al. (2023); Wu et al. (2023); Ji et al. (2023) train LLMs on varying combinations of such preferential properties. In practice, one tends to resort to trial and error to find the right preference combination for their particular task. In doing so, one verifies if certain minimum baselines are satisfied, such as ensuring the factual correctness of statements or confirming that the response lengths are capped at 100 words. Since there isn’t a way to enforce such requirements directly, current methods resort to heuristics. Additionally, existing pipelines carry out SFT and alignment in sequential steps. Hence they need to ensure that the LLM does not forget relevant task information learned during the SFT stage. This is achieved by penalizing the LLM for drastic deviations: the strength of the penalty is determined heuristically.
@@
In this work, we formulate SFT and alignment in LLMs as a constrained optimization problem. In particular, we fine-tune an LLM to minimize the task perplexity (the objective function) while simultaneously satisfying application-specific minimum requirements (the constraints). This merges the SFT and alignment stages and mitigates the reliance on heuristics altogether. Furthermore, we propose Lagrange Large Language Models, a.k.a. L3Ms, to solve such constrained optimization problems. Specifically, they do so by employing logarithmic barriers and enforcing the constraints gradually over the training procedure. Lastly, we empirically demonstrate how one can pick and choose different constraints and tailor L3Ms to a range of applications without resorting to heuristics. For example, Table 1 provides the generated responses from two such L3Ms; both are trained on the task of instruction-following, but one is constrained to be concise while the other is to be verbose.
@@
    We demonstrate experimentally how L3Ms can be customized to different applications and their specific requirements while avoiding heuristic-driven processes (cf. Section 6).
@@
Table 1: Example responses from length constrained L3Ms. We provide example responses from L3Ms with varying length constraints. We include the prompt along with the generated responses from two L3Ms: one constrained to have short responses, and the other constrained to long ones.
@@ 2.1 Pre-training
where $x_{i}$ and $x_{<i}$ denote the $i$-th token in the sequence and its prefix, respectively. Additionally, the function $\pi_{\theta}(\cdot)$ denotes the LLM’s predicted probability distribution over token sequences, where the LLM is parameterized with weights $\theta$. Then, the pre-training objective is given as follows:
@@ 2.2 Supervised fine-tuning (SFT)
Next, one fine-tunes the LLM to a task of interest, such as instruction-following, summarization, or translation. The data are (prompt, response) pairs $(x,y)$, and the LLM is trained to regenerate these responses. Thus, one minimizes the perplexity on the response conditioned on the prompt:
@@
for the distribution $p(\cdot)$ over (prompt, response) pairs which reflects the task-related data.
@@ 2.3 Alignment
This stage aligns the LLM to generate responses with preferential properties. A common setup is to learn preference reward functions that represent properties like helpfulness and harmlessness (Bai et al., 2022), followed by reinforcement learning to adapt the LLM to maximize said rewards. This is referred to as reinforcement learning from human feedback (RLHF; Knox & Stone, 2008; Griffith et al., 2013; Christiano et al., 2017). Note that the preference reward functions need not always be learned; they could also be engineered or rule-based, such as the length of the response.
Given a single preference reward function $r(\cdot)$, the alignment objective is given as follows:
@@
This maximizes the rewards for the LLM’s responses to prompts sampled from the task distribution. To prevent over-optimization of the reward and avoid drastic deviation away from the SFT model, it is common practice to add a regularization penalty such as the KL divergence (Gao et al., 2023).
We are interested in $k\geq 1$ different preferential properties. In such a scenario, one could learn individual preference reward functions $r_{i}$’s and optimize the LLM to maximize their combination. In particular, Li et al. (2021); Rame et al. (2023); Wu et al. (2023); Ji et al. (2023) use a linear combination of the rewards, substituting the single reward in Eq. 2 with $\sum_{i=1}^{k}\alpha_{i}r_{i}(y|x)$, for some choice of $\alpha_{i}\geq 0$. Alternatively, one could learn a single reward function by choosing the data proportions of the different properties used to train it. For instance, Bai et al. (2022) use a 3:1 proportion of helpfulness to harmlessness data to train a single preference reward function. Note that the proportions can also be represented via weights $\alpha_{i}\geq 0$ with $\sum_{i=1}^{k}\alpha_{i}=1$. Therefore, the choice of $\alpha_{i}$’s, combined with the strength of the regularization penalty, together steer the alignment. They are chosen based on the good judgment of the practitioner in a somewhat heuristic manner.
@@ 2.4 Shortcomings
While the above pipeline is commonly used to train LLMs for deployment, it has shortcomings.
@@
Secondly, recall that the original task objective is in Eq. 1. However, optimizing a special purpose objective like Eq. 2 can lead to a decrease in performance on the original task. While penalizing the deviation away from the SFT model helps mitigate this to some extent, the strength of the penalty is again determined heuristically. Can we ensure the performance on the original task is maintained?
@@ 3.1 One size does not fit all
Early work on LLM alignment assumed homogeneity of preferences (Bakker et al., 2022). However, the reality is quite different: human preferences vary widely and are highly application-specific (Rame et al., 2023; Casper et al., 2023). Consequently, Rame et al. (2023); Wu et al. (2023); Ji et al. (2023); Zhou et al. (2024) linearly combine several fine-grained preference rewards, where each combination reflects a different customized preference. However, the right combination of preferential properties for a particular application is determined heuristically through trial and error. Conversely, we formulate alignment as a constrained optimization problem, where the custom requirements are naturally imposed as constraints. This alleviates the need for heuristic choices.
@@ 3.2 Constrained alignment for LLMs
Independent of our work, Moskovitz et al. (2024); Dai et al. (2024) also introduced constrained optimization for LLM alignment but for varying reasons. Motivated to avoid over-optimization of preference rewards, Moskovitz et al. (2024) find “proxy points”, values of the reward functions beyond which the performance of the LLM is negatively impacted. They constrain the average rewards to be in the vicinity of these proxy points. Dai et al. (2024) trade-off between helpfulness and harmlessness from a safety standpoint. They maximize the LLM’s helpfulness while constraining its average harmlessness reward. Our motivation on the other hand is to tailor to custom preferences through sets of different constraints, while simultaneously learning to perform the task at hand.
Our work is different in two important ways. Firstly, both Moskovitz et al. (2024); Dai et al. (2024) consider the alignment process in isolation, because of which their objective is either to maximize one of the preference rewards or to minimize the deviation away from the SFT model. On the other hand, we merge the SFT and alignment stages and choose task perplexity minimization as our objective. This directly ensures that the LLM’s task-solving capabilities are preserved, and it avoids loading an additional reference (SFT) model during the training process to compute the deviations.
Secondly, both Moskovitz et al. (2024); Dai et al. (2024) obtain solutions to the constrained problem by computing the saddle-point of the Lagrangian. This is done by formulating a min-max game where the LLM minimizes the Lagrangian and the Lagrange multipliers adapt to maximize it. However, the non-convexity of the objective makes this nontrivial. Instead, we choose the logarithmic barrier approach as the associated Lagrange multipliers satisfy the KKT complementary slackness condition by design (cf. Section 5.2) rather than letting the learning procedure do so (which in our experience is extremely sensitive to the choice of the learning rate). We empirically observe that while both methods satisfy the constraints, our approach using log-barriers minimizes the task perplexity better (cf. Section 6.3). As a side effect, we avoid introducing new learning parameters.
Additionally, we provide experimental validation of our approach on more preferences than Moskovitz et al. (2024); Dai et al. (2024), and a 4.5$\times$ larger LLM than Moskovitz et al. (2024).
4 Constrained optimization
Our goal is to reduce the task objective in Eq. 1 for the LLM to solve the intended task, while also enabling custom alignment by having the LLM meet the application-specific minimum requirements on different preference rewards. To do so, we propose the constrained optimization problem:
@@ 4 Constrained optimization
Here, the objective is the same as the task objective in SFT, and the constraints are enforced to satisfy the custom requirements; this merges the SFT and alignment stages. The $b_{i}$’s signify the minimum baselines for each preference reward function $r_{i}(\cdot)$ and the constraints are enforced on average.
Compared to the previous approach, we no longer rely on heuristics to find a set of weights $\alpha_{i}$’s to satisfy the minimum requirements; we can do so directly through the constraints. Furthermore, with the same objective as SFT, we can directly maintain task performance without any deviation penalty. Additionally, note that whenever a constraint is satisfied, its influence vanishes. For instance, if the LLM is naturally harmless and $r_{\text{harmless}}(y|x)\geq b_{\text{harmless}}$, then the constraint is not active and the LLM will not be penalized. Conversely, the previous approach would further penalize the LLM.
@@
For ease of notation, we rewrite the constrained problem in Eqs. 3 and 4 as follows:
@@ 4.1 Types of constraints
While these constraints are not equivalent, they are related. We can rewrite Eq. 7 in the form of expectation constraints by using $1-\epsilon_{i}$ as the threshold and taking the expectation of the indicator $\mathds{1}\{r_{i}(y|x)\geq b_{i}\}$. Moreover, Eq. 6 implies Eq. 4, but the converse is not true. Unfortunately, Eq. 6 is difficult to achieve in practice, especially when the data distribution is unknown.
We continue using the expectation constraints, but similar discussions can extend to the other types.
@@ 4.2 Lagrange multipliers
There is a rich literature connecting the Lagrangian with constrained optimization. Notably, the KKT conditions (Karush, 1939; Kuhn & Tucker, 1951) provide sufficiency conditions for global optimality under convexity, where the solution can be obtained by finding the saddle point of the Lagrangian. However, these conditions are not enough for highly non-convex scenarios like ours.
Nonetheless, the Lagrangian is instructive in understanding the relative importance of the constraints. For an active constraint, i.e., one satisfied with equality, the corresponding Lagrange multiplier can be non-zero; the larger its value, the more important the constraint. Conversely, for an inactive constraint, i.e., one satisfied with strict inequality, the corresponding Lagrange multiplier must vanish to 0. This is known as complementary slackness and is one of the KKT conditions.
@@ 4.3 Logarithmic barrier
A practical way to enforce constraints is with barrier functions. Consider the (relaxed) log-barrier:
@@
with parameters $\mu,s>0$. This is a convex, continuous, and differentiable function, which is valid for all $z\in\mathds{R}$. Importantly, for $s=\mu^{2}$, this barrier function converges to the characteristic function $\chi\{z\leq 0\}$ as $\mu\rightarrow 0$, i.e., it takes the value 0 when $z\leq 0$ and $\infty$ otherwise (Tal et al., 1992; Nash et al., 1994; Hauser & Saccon, 2006; Feller & Ebenbauer, 2017); the condition $s=\mu^{2}$ is sufficient, but not necessary (Kervadec et al., 2022). This convergence to the characteristic function is visually depicted in Fig. 1, showing the change in the log-barrier function as we gradually decrease $\mu$.
We can now use the log-barrier to enforce the constraints in Eq. 5 and simply add them to the objective. We obtain an unconstrained objective, with $\mu$ controlling the strength of the constraints:

### v2

@@ Abstract
Supervised fine-tuning (SFT) and alignment of large language models (LLMs) are key steps in providing a good user experience. However, the concept of an appropriate alignment is inherently application-dependent, and current methods often rely on heuristic choices to drive optimization. In this work, we formulate SFT and alignment as a constrained optimization problem: the LLM is fine-tuned on a task while being required to meet application-specific requirements, without resorting to heuristics. To solve this, we propose Lagrange Large Language Models (L3Ms), which employ logarithmic barriers to enforce the constraints. This approach allows for the customization of L3Ms across diverse applications while avoiding heuristic-driven processes. We experimentally demonstrate the versatility and efficacy of L3Ms in achieving tailored alignments for various applications.
@@ 1 Introduction
Large language models (LLMs) are used for a wide range of tasks: as chatbots (Brown et al., 2020; OpenAI, 2024), for code generation (Ahmad et al., 2021; Wang et al., 2021; Rozière et al., 2024), for medical assistance (Yang et al., 2022; Moor et al., 2023), and so on. The key ingredients for their impressive downstream performance are supervised fine-tuning (SFT) and alignment; the former fine-tunes the LLM to a task of interest, while the latter instills it with preferential properties. Arguably, the right combination of preferential properties is highly application-dependent. For example, a scholar would want a chatbot to be honest and factual for assistance with their work. In contrast, a fiction writer might prefer the opposite behavior to help create fantastical imaginary worlds. There is also plenty of (anecdotal) evidence in support: some LLMs refuse to provide information on how to “kill” a process in Unix, recommending the use of less violent strategies for dealing with wayward computer programs instead. [^1] https://www.reddit.com/r/LocalLLaMA/comments/180p17f/new_claude_21_refuses_to_kill_a_python_process/ Therefore, we need frameworks for LLM customization.
@@
Consequently, Bai et al. (2022); Rame et al. (2023); Wu et al. (2023); Ji et al. (2023); Zhou et al. (2024) fine-tune LLMs on varying combinations of such preferential properties. In practice, one tends to resort to trial and error to find the right combination of preferences for their particular application. In doing so, one verifies if certain minimum baselines are satisfied, such as ensuring the factual correctness of statements or confirming that response lengths are capped at 100 words. Since there isn’t a way to enforce such requirements directly, current methods resort to heuristics. Additionally, existing pipelines carry out SFT and alignment sequentially and must ensure that the LLM does not forget relevant task information learned during the SFT stage. This is achieved by penalizing the LLM for drastic deviations, with the strength of the penalty determined heuristically.
@@
In this work, we formulate SFT and alignment in LLMs as a constrained optimization problem. In particular, we fine-tune an LLM to minimize the task objective (the objective function) while simultaneously satisfying application-specific minimum requirements (the constraints). This merges the SFT and alignment stages and mitigates the reliance on heuristics altogether. Furthermore, we propose Lagrange Large Language Models, a.k.a. L3Ms, to solve such constrained optimization problems. Specifically, we do so by employing logarithmic barriers and gradually enforcing the constraints during the fine-tuning procedure. Lastly, we empirically demonstrate how one can pick and choose constraints and tailor L3Ms to a range of applications without resorting to heuristics. For example, Table 1 provides the generated responses from two such L3Ms; both are fine-tuned for instruction-following, but one is constrained to be concise, while the other is to be verbose.
@@
    We experimentally demonstrate how L3Ms can be customized to different applications and their specific requirements while avoiding heuristic-driven processes (cf. Section 6).
@@
Table 1: Example responses from length constrained L3Ms. We provide example responses from L3Ms with varying length constraints. We include the prompt along with the generated responses from two L3Ms; one constrained to have short responses and the other constrained to long ones.
@@ 2.1 Pre-training
where $x_{i}$ and $x_{<i}$ denote the $i$-th token in the sequence and its prefix, respectively. Additionally, the function $\pi_{\theta}(\cdot)$ denotes the predicted probability distribution of the LLM over token sequences, where the LLM is parameterized with weights $\theta$. Then, the pre-training objective is given as:
@@ 2.2 Supervised fine-tuning (SFT)
Next, one fine-tunes the LLM to a task of interest, such as instruction-following, summarization, or translation. The data are (prompt, response) pairs $(x,y)$, and the LLM is fine-tuned to regenerate the responses. Thus, one minimizes the perplexity (or a related loss) on the response given the prompt:
@@
for a distribution $p(\cdot)$ over (prompt, response) pairs that reflects the task-related data.
@@ 2.3 Alignment
This stage aligns the LLM to generate responses with preferential properties. A common setup is to learn preference reward functions that represent properties such as helpfulness and harmlessness (Bai et al., 2022), followed by reinforcement learning to adapt the LLM to maximize the said rewards. This is called reinforcement learning from human feedback (RLHF; Knox & Stone, 2008; Griffith et al., 2013; Christiano et al., 2017). Note that the preference reward functions need not always be learned; they could also be engineered or rule-based, such as the length of the response.
Given a single preference reward function $r(\cdot)$, the alignment objective is given as:
@@
This maximizes the rewards for the LLM’s responses to prompts sampled from the task distribution. To prevent over-optimization of the reward and avoid drastic deviation away from the SFT model, it is common practice to add a regularization penalty, such as the KL divergence (Gao et al., 2023).
We are interested in $k\geq 1$ different preferential properties. In such a scenario, one could learn individual preference reward functions $r_{i}(\cdot)$’s and optimize the LLM to maximize their combination. In particular, Li et al. (2021); Rame et al. (2023); Wu et al. (2023); Ji et al. (2023) use a linear combination of the rewards, substituting the single reward in Eq. 2 with $\sum_{i=1}^{k}\alpha_{i}r_{i}(y|x)$, for some choice of $\alpha_{i}\geq 0$. Alternatively, one could learn a single reward function by choosing the data proportions of the different properties used to train it. For example, Bai et al. (2022) use a 3:1 proportion of helpfulness to harmlessness data to train a single preference reward function. Note that the proportions can also be represented as weights $\alpha_{i}\geq 0$ with $\sum_{i=1}^{k}\alpha_{i}=1$. Therefore, the choice of $\alpha_{i}$’s, combined with the strength of the regularization penalty, together steer the alignment. They are chosen based on the good judgment of the practitioner in a somewhat heuristic manner.
@@ 2.4 Shortcomings
Although the above pipeline is commonly used to train LLMs for deployment, it has shortcomings.
@@
Secondly, recall that the original task objective is given in Eq. 1. However, optimizing a special purpose objective such as Eq. 2 can lead to a decrease in performance on the original task. Whilst penalizing the deviation away from the SFT model mitigates this to some extent, the strength of the penalty is again determined heuristically. Can we ensure the performance on the original task?
@@ 3.1 One size does not fit all
Early work on LLM alignment assumed homogeneity of preferences (Bakker et al., 2022). However, the reality is quite different: human preferences vary widely and are highly application-specific (Rame et al., 2023; Casper et al., 2023). Consequently, Bai et al. (2022); Rame et al. (2023); Wu et al. (2023); Ji et al. (2023); Zhou et al. (2024) linearly combine several fine-grained preference rewards, where each combination reflects a unique customization. However, the right combination of preferential properties for a particular application is determined heuristically by trial and error. In contrast, we formulate alignment as a constrained optimization problem. In doing so, custom requirements are naturally imposed as constraints. This alleviates the need for heuristic choices.
@@ 3.2 Constrained alignment for LLMs
Independent of our work, Moskovitz et al. (2024); Dai et al. (2024) also introduced constrained optimization for LLM alignment but for varying reasons. Motivated to avoid over-optimization of preference rewards, Moskovitz et al. (2024) find “proxy points”, values of the reward functions beyond which the performance of the LLM is negatively impacted. They constrain the average rewards to be in the vicinity of these proxy points. Dai et al. (2024) trade-off between helpfulness and harmlessness from a safety perspective. They maximize the LLM’s helpfulness while constraining its average harmlessness reward. Our motivation, on the other hand, is to tailor to custom preferences through sets of different constraints while simultaneously learning to perform the task of interest.
Our work is different in two important ways. Firstly, both Moskovitz et al. (2024); Dai et al. (2024) consider the alignment process in isolation, because of which their objective is either to maximize one of the preference rewards or to minimize the deviation away from the SFT model. Instead, we merge the SFT and alignment stages and minimize the task objective directly. This ensures that the LLM learns task-solving capabilities without any deviation penalty. Furthermore, by avoiding the computation of the deviations from the SFT model, we save both memory and time during the fine-tuning process: we no longer need to load the SFT model and evaluate it on generated responses.
Secondly, both Moskovitz et al. (2024); Dai et al. (2024) obtain a solution to the constrained problem by computing a saddle point of the Lagrangian. This is achieved by formulating a minimax game where the LLM minimizes the Lagrangian, and the Lagrange multipliers adapt to maximize it. However, the non-convexity of the objective makes this nontrivial. Instead, we choose the logarithmic barrier approach as the associated Lagrange multipliers satisfy the KKT complementary slackness condition by design (cf. Section 5.2) rather than letting the learning procedure do so (which in our experience is extremely sensitive to the choice of the learning rate). We empirically observe that while both methods satisfy the constraints, our approach using log barriers minimizes the task objective better (cf. Section 6.3). As a side effect, we avoid introducing new learning parameters.
Furthermore, we provide experimental validation of our approach using more preferences than Moskovitz et al. (2024); Dai et al. (2024) and 4.5$\times$ larger LLMs than Moskovitz et al. (2024).
4 Constrained optimization for LLMs
Our goal is to reduce the task objective for the LLM to solve the task of interest while also enabling custom alignment by having the LLM meet the application-specific minimum requirements on different preference rewards. To do so, we propose the constrained optimization problem:
@@ 4 Constrained optimization for LLMs
Here, the objective is the same as that of SFT in Eq. 1, and the constraints are enforced to satisfy the custom requirements; this merges the SFT and alignment stages. The $b_{i}$’s signify the minimum baselines for each preference reward function $r_{i}(\cdot)$, and the constraints are enforced on average.
Compared to the previous approach, we no longer rely on heuristics to find a set of weights $\alpha_{i}$’s to satisfy the minimum requirements; we can do so directly through the constraints. Furthermore, with the same objective as SFT, we can directly maintain task performance without any deviation penalty. Additionally, note that whenever a constraint is satisfied, its influence vanishes. For example, if the LLM is naturally harmless and $r_{\text{harmless}}(y|x)>b_{\text{harmless}}$, then the constraint is not active, and the LLM will not be penalized. In contrast, the previous approach would further penalize the LLM.
@@
For ease of notation, we rewrite the constrained problem in Eqs. 3 and 4 as:
@@ 4.1 Types of constraints
These constraints are not equivalent, but they are related. We can rewrite Eq. 7 in the form of average constraints using $1-\epsilon_{i}$ as the threshold and taking the expectation of the indicator $\mathds{1}\{r_{i}(y|x)\geq b_{i}\}$. Moreover, Eq. 6 implies Eq. 4, but the converse is not true. Unfortunately, Eq. 6 is difficult to achieve in practice, especially when the data distribution is unknown.
We continue using expectation/average constraints, but similar discussions can extend to other types.
@@ 4.2 Lagrange multipliers
There is a rich literature connecting the Lagrangian with constrained optimization. Notably, the KKT conditions (Karush, 1939; Kuhn & Tucker, 1951) provide sufficiency conditions for global optimality under convexity, where the solution is obtained by finding the saddle point of the Lagrangian. However, these conditions are not enough for highly non-convex scenarios such as ours.
Nevertheless, the Lagrangian is instructive in understanding the relative importance of the constraints. For an active constraint, i.e., one satisfied with equality, the corresponding Lagrange multiplier can be non-zero; the larger its value, the more important the constraint. Conversely, for an inactive constraint, i.e., one satisfied with strict inequality, the corresponding Lagrange multiplier must vanish to 0. This is known as complementary slackness and is one of the KKT conditions.
@@ 4.3 Logarithmic barrier
A practical way to enforce constraints is with barrier functions. Consider the (relaxed) log barrier:
@@
with parameters $\mu,s>0$. This is a convex, continuous, and differentiable function, which is valid for all $z\in\mathds{R}$. Importantly, for $s=\mu^{2}$, this barrier function converges to the characteristic function $\chi\{z\leq 0\}$ as $\mu\rightarrow 0$, i.e., it takes the value 0 when $z\leq 0$ and $\infty$ otherwise (Tal et al., 1992; Nash et al., 1994; Hauser & Saccon, 2006; Feller & Ebenbauer, 2017); the condition $s=\mu^{2}$ is sufficient, but not necessary (Kervadec et al., 2022). This convergence to the characteristic function is visually depicted in Fig. 1, showing the change in the log barrier function as we gradually decrease $\mu$.
We can now use the log barrier to enforce the constraints in Eq. 5 and simply add them to the objective. We obtain an unconstrained objective, with $\mu$ controlling the strength of the constraints: